OpenStudy (anonymous):

A drinking glass is shaped like a cylinder with a height of 12 cm and a radius of 4 cm. Maeva adds 25 spherical pieces of ice to the glass. The pieces of ice each have a diameter of 3 cm. How many cubic centimeters of water can Maeva add to fill the glass to the rim? Use 3.14 to approximate pi and express your answer in hundredths. ___cm3

OpenStudy (anonymous):

@lsherron

OpenStudy (anonymous):

I need help in like maybe 5 mins

OpenStudy (anonymous):

I got 150.72 by doing the math for it, but I am at least 99% sure

OpenStudy (anonymous):

Did u multiply pi too?

OpenStudy (anonymous):

@lsherron

OpenStudy (anonymous):

that is the anser

OpenStudy (anonymous):

I dunno I need two more before submitting

OpenStudy (anonymous):

A non-scale model of the earth has a diameter of 18 in. Within an outer clay layer, there is a core made of a ball with a diameter of 3 in. What is the volume of the clay? Use 3.14 to approximate pi and express your answer in hundredths. ___in3

OpenStudy (anonymous):

Fist find volume of whole find volue of golf ball volume of sphere=4/3 times pi times r^3 diameter=2radius divide 2 diameter/2=radius 12/2=6=radius 4/3 times 3.14 times 6^3=904.32 whole volume=904.32 volume of golve ball 4/3 times 3.14 times 2/2=4.18666 subtract 904.32-4.18666=900.133333 the asnwer is 900.13 in^3

OpenStudy (anonymous):

3037.95 is the right answer

OpenStudy (anonymous):

my brother did that

OpenStudy (anonymous):

Ok thx

OpenStudy (anonymous):

A can of spaghetti has a diameter of 12 cm and a height of 16 cm. The can contains spherical meatballs, each with a radius of 3 cm, that take up 25% of the volume. How many meatballs are in the can? Use 3.14 to approximate pi and express your answer as a whole number.

OpenStudy (anonymous):

Raduis of Can=1/2 of 12cm R=6cm Volume of Can= pi r ^{2} h = pi 6 ^{2} 16 =1809.5573cm3 Volume of Meatball=4/3 pi r ^{3} =4/3 pi 3 ^{3} =113.09734 25% of can = 1809.5573/4 =452.389325 Meatballs in area = 25%of can/volume of meatballs =452.389325/113.09734 =4 meatballs

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

np any more

OpenStudy (anonymous):

No the first one was 249.63 and the last two were right thx

OpenStudy (anonymous):

what was it the grade

OpenStudy (anonymous):

40%

OpenStudy (anonymous):

aw man im sorry